Grade 11General Physics

The speed of a man is 4m/s,the speed of river is 2m/s.Find the angle between the man and river so that he travel in a straight line perpendicular to the river.

Seerat
7 Years agoGrade 11

1 Answer

Piyush Kumar Behera
7 Years ago
Let us understand how should man travel so that it reaches straight.
If man tends to move to left of the river then there is a possibility that the river drags back to the same vertical position and makes him to reach at the other side of the bank.
So Vrelative must have a left horizontal component whose magnitude must be equal to the river velocity so that there is no net displacement along the river.
So Vhorizontal = 2m/s
Total velocity = 4m/s
vertical component=\sqrt{4^{2}-2^{2}}} = 2\sqrt{3} m/s
 
So angle b/w man and river is 90+tan^{-1}{\frac{1}{\sqrt{3}}} = 120 degrees
 
 
So the answer is 120 degree.
I hope it helps!!
 
Piyush Kumar Behera
IIT Dhanbad